5. A spring exerts a force of 15 N to propel a 125 g toy car forward along a table top. a) If the toy car was initially motionless and the force was applied over a distance of 12 cm, what is the theoretical maximum speed of the toy? /2 b) If the toy car then travels up a ramp, how high will it travel vertically before coming to a complete stop? 12 6. A tov cart of mass 5kg is projected up a ramp inclined at 30° to the berizente
1
Expert's answer
2020-10-29T07:01:11-0400

a) According to Newton's second law, the acceleration is


"a=\frac{F}{m}."

The speed after uniformly accelerated motion across distance d is


"v=\sqrt{2ad}=\sqrt{\frac{2Fd}{m}}=5.4\text{ m/s}."

b) Apply conservation of energy principle: all kinetic energy converts to potential when the car stops.


"mgh=\frac{mv^2}2,\\
h=\frac{v^2}{2g}=1.5\text{ m}."